given the function h(x)=4x+4 find each of the following
h(4)=
h(-10)=
h(6,6)=

Answer by subudear(62)

Given h(x)=4x+4
h(4)= 4*4+4 = 20
h(-10)= 4*(-10)+4 = -36
h(6,6)= 4*6+4 = 28